# Energy-Based Model for Accurate Estimation of Shapley Values in Feature Attribution

### v3

@@ VII-B Error Bound of VAEAC
$=\mathbb{E}_{q(\mathbf{z}|\mathbf{x})}\log p(\mathbf{x})$
$=\mathbb{E}_{q(\mathbf{z}|\mathbf{x})}\log\frac{p(\mathbf{x},\mathbf{z})}{p(\mathbf{z}|\mathbf{x})}$
$=\mathbb{E}_{q(\mathbf{z}|\mathbf{x})}\log\frac{p(\mathbf{x},\mathbf{z})q(\mathbf{z}|\mathbf{x})}{p(\mathbf{z}|\mathbf{x})q(\mathbf{z}|\mathbf{x})}$
$=\mathbb{E}_{q(\mathbf{z}|\mathbf{x})}\log\frac{p(\mathbf{x},\mathbf{z})}{q(\mathbf{z}|\mathbf{x})}+\mathbb{E}_{q(\mathbf{z}|\mathbf{x})}\log\frac{q(\mathbf{z}|\mathbf{x})}{p(\mathbf{z}|\mathbf{x})}$
$=\mathbb{E}_{q(\mathbf{z}|\mathbf{x})}\log\frac{p(\mathbf{x},\mathbf{z})}{q(\mathbf{z}|\mathbf{x})}+D_{KL}\left(q(\mathbf{z}|\mathbf{x})||p(\mathbf{z}|\mathbf{x})\right),$
where $\mathbb{E}_{q(\mathbf{z}|\mathbf{x})}\log\frac{p(\mathbf{x},\mathbf{z})}{q(\mathbf{z}|\mathbf{x})}$ is called evidence lower bound (ELBO) and $D_{KL}(\cdot||\cdot)$ is Kullback-Leibler (KL) divergence. ELBO is the objective function that needs to be optimized, by maximizing ELBO via gradient descent, the parameters of VAE can be obtained.
VAEAC, a Shapley value approximation approach based on VAE, is introduced by Ref. [17]. In VAEAC, the contribution function of Shapley value is sampled from the conditional probability $p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ generated by VAE. The log-likelihood introduced in Ref. [17] is as follows.
$\quad\log p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ (B4)
$=\mathbb{E}_{q(\mathbf{z}|\mathbf{x}_{\bar{S}},\mathbf{x}_{S})}\left[\log\frac{p(\mathbf{x}_{\bar{S}},\mathbf{z}|\mathbf{x}_{S})}{q(\mathbf{z}|\mathbf{x}_{\bar{S}},\mathbf{x}_{S})}\right]+\mathbb{E}_{q(\mathbf{z}|\mathbf{x}_{\bar{S}},\mathbf{x}_{S})}\left[\log\frac{q(\mathbf{z}|\mathbf{x}_{\bar{S}},\mathbf{x}_{S})}{p(\mathbf{z}|\mathbf{x}_{\bar{S}},\mathbf{x}_{S})}\right]$
$=\mathbb{E}_{q(\mathbf{z}|\mathbf{x}_{\bar{S}},\mathbf{x}_{S})}\left[\log\frac{p(\mathbf{x}_{\bar{S}},\mathbf{z}|\mathbf{x}_{S})}{q(\mathbf{z}|\mathbf{x}_{\bar{S}},\mathbf{x}_{S})}\right]+D_{KL}(q(\mathbf{z}|\mathbf{x}_{\bar{S}},\mathbf{x}_{S})||p(\mathbf{z}|\mathbf{x}_{\bar{S}},\mathbf{x}_{S}))$
$\geq\mathbb{E}_{q(\mathbf{z}|\mathbf{x}_{\bar{S}},\mathbf{x}_{S})}\left[\log\frac{p(\mathbf{x}_{\bar{S}},\mathbf{z}|\mathbf{x}_{S})}{q(\mathbf{z}|\mathbf{x}_{\bar{S}},\mathbf{x}_{S})}\right].$
VAEAC trains the parameters by maximizing the last term in Eq.(B4), and then samples from the generated conditional distribution $p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ to calculate the contribution function $\hat{v}(S)$.
According to Ref. [18], the inference error of VAE can be decomposed into two components, the approximation error and the amortization error. Since the analysis of the model error bound in this paper assumes the model can achieve theoretical optimal performance after a large number of iterative training, the approximation error of VAEAC, i.e. $\log p_{data}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})-p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$, can be controlled and upper bounded by a small positive number. Thus, similar to Assumption 1, the probability density difference between $p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})$ and $p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})$ is bounded as follows.
$\small\left|\frac{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}-1\right|\leq\epsilon_{1},$ (B5)
Another non-negligible error of VAEAC is the amortization error, i.e. $\log p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})-\log p_{\theta}(\mathbf{x}_{\bar{S}|\mathbf{x}_{S}})$, which arises because VAE uses a shared encoder to approximate the posterior distribution for all data points, rather than optimizing a distinct posterior distribution for each individual sample. This amortization error also leads to the following difference between the probability densities.
$\small\left|\frac{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}-1\right|\leq\delta.$ (B6)
$\mathbb{E}[|\hat{v}(S)-v(S)|]$ of VAEAC can be derived as follows.
$\quad\mathbb{E}[|\hat{v}(S)-v(S)|]$ (B7)
$=\mathbb{E}\left[\left|\frac{1}{K}\sum_{k=1}^{K}f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})-\mathbb{E}_{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right|\right]$
$=\mathbb{E}\left[\left|\frac{1}{K}\sum_{k=1}^{K}f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})-\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right.\right.$
$\quad\left.\left.+\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]-\mathbb{E}_{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right.\right.$
$\quad\left.\left.+\mathbb{E}_{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]-\mathbb{E}_{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right|\right]$
$\quad+\mathbb{E}\left[\left|\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]-\mathbb{E}_{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right.\right.$
$\quad\left.\left.+\mathbb{E}_{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]-\mathbb{E}_{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}^{t}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right|\right].$
The first term in the last inequality of Eq.(B7) is the statistical error, the second and the third terms are related to the amortization error and approximation error of VAE, respectively. The upper bound of the statistical error is as follows.
$\small\mathbb{E}\left[\left|\frac{1}{K}\sum_{k=1}^{K}f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})-\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right|\right]\leq\frac{\sqrt{\pi}}{\sqrt{2K}}.$ (B8)
The amortization error is bounded as follows.
$\quad\mathbb{E}\left[\left|\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]-\mathbb{E}_{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right|\right]$ (B9)
$=\mathbb{E}\left[\left|\int f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})d\mathbf{x}_{\bar{S}}-\int f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})d\mathbf{x}_{\bar{S}}\right|\right]$
$=\mathbb{E}\left[\left|\int f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\frac{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})d\mathbf{x}_{\bar{S}}\right.\right.$
$\left.\left.\quad-\int f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})d\mathbf{x}_{\bar{S}}\right|\right]$
$=\mathbb{E}\left[\left|\mathbb{E}_{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[\left(f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\frac{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}-f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})\right)\right]\right|\right]$
$\leq\mathbb{E}\left[\mathbb{E}_{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[\left|f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\frac{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}-f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})\right|\right]\right]$
$=\delta.$
The approximation error is bounded as follows,
$\quad\mathbb{E}\left[\left|\mathbb{E}_{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]-\mathbb{E}_{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right|\right]$ (B10)
$=\mathbb{E}\left[\left|\int f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})d\mathbf{x}_{\bar{S}}-\int f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})d\mathbf{x}_{\bar{S}}\right|\right]$
$=\mathbb{E}\left[\left|\int f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\frac{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})d\mathbf{x}_{\bar{S}}\right.\right.$
$\quad-\left.\left.\int f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})d\mathbf{x}_{\bar{S}}\right|\right]$
$=\mathbb{E}\left[\left|\mathbb{E}_{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[\left(f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\frac{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}-f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})\right)\right]\right|\right]$
$\leq\mathbb{E}\left[\mathbb{E}_{p_{data}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[\left|f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\frac{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}-f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})\right|\right]\right]$
$\leq\mathbb{E}\left[\mathbb{E}_{p_{data}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[\left|\left(\frac{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}-1\right)f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})\right.\right.\right.$
$\quad\left.\left.\left.+f^{\prime}(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})(\mathbf{x}_{S}-\mathbf{x}_{S}^{t})+o(\mathbf{x}_{S}-\mathbf{x}_{S}^{t})\right|\right]\right]$
$\leq\epsilon_{1}f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})+\epsilon_{1}^{\prime}$
$\leq\epsilon_{2}.$
By combining Eqs. (B8), (B9), and (B10), the error bound of VAEAC approach can be obtained.
$\small\mathbb{E}[|\hat{v}(S)-v(S)|]=\frac{\sqrt{\pi}}{\sqrt{2K}}+\epsilon_{2}+\delta.$ (B11)
It can be seen from Eq.(B11), comparing with EmSHAP, the error bound of VAEAC contains the term $\delta$ that is related to the amortization error.

### v4

@@ VII-F Performance analysis of EmSHAP under different parameters and model structures
ADTI
[…]
ETT
[…]
MIMIC
[…]
TABLE E2: SIC AUC results for different context vector dimensions
Dataset
SIC AUC
Dimension of context vector
0 8 16 32 64 128
MNIST
[…]
ADTI
[…]
ETT
[…]
MIMIC
[…]
TABLE E3: Performance of EmSHAP under different latent dimensions
Dataset
Latent
dimension
SIC AUC Run time Memory usage Energy consumption
[…]
MNIST
16 0.931 0.152 2813.42s 1.46s 1706.36MB 1788.95MB 7.23KJ 1.49KJ
32 0.965 0.037 2977.10s 1.63s 1823.91MB 1879.68MB 8.62KJ 2.13KJ
64 0.965 0.033 3460.16s 4.63s 1946.19MB 2024.79MB 9.48KJ 2.63KJ
128 0.966 0.031 4103.57s 13.89s 2237.85MB 2242.40MB 11.21KJ 3.15KJ
ADTI
16 0.921 0.466 230.49s 0.33s 918.45MB 943.34MB 11.38KJ 0.05KJ
32 0.933 0.411 270.90s 0.87s 966.58MB 946.81MB 15.53KJ 0.05KJ
64 0.950 0.368 303.83s 1.04s 1019.57MB 1046.63MB 19.12KJ 0.07KJ
128 0.951 0.355 326.36s 1.84s 1120.41MB 1148.21MB 22.45KJ 0.09KJ
ETT
16 0.884 0.256 10.76s 2E-3s 519.68MB 538.77MB 0.82KJ 0.03KJ
32 0.915 0.243 16.55s 4E-3s 554.11MB 540.46MB 0.82KJ 0.03KJ
64 0.920 0.227 25.45s 0.10s 611.91MB 638.35MB 0.82KJ 0.03KJ
128 0.923 0.220 32.17s 0.39s 649.00MB 653.30MB 0.83KJ 0.03KJ
MIMIC
16 0.600 0.393 851.10s 58.15s 1159.61MB 1169.85MB 27.71KJ 8.81KJ
32 0.699 0.384 1168.72s 59.93s 1161.51MB 1168.10MB 28.03KJ 9.16KJ
64 0.700 0.375 1394.08s 72.99s 1261.87MB 1217.31MB 29.35KJ 9.37KJ
128 0.710 0.372 2266.46s 86.59s 1367.76MB 1320.65MB 53.79KJ 10.33KJ
TABLE E4: Comparison of different RNN-based proposal networks
Dataset
Proposal
network
SIC AUC Run time Memory usage
[…]
MNIST
LSTM 0.952 0.108 3406.11s 2.01s 2067.50MB 1999.63MB
GRU 0.965 0.037 2977.10s 1.63s 1879.91MB 1823.68MB
BiGRU 0.942 0.084 4108.33s 2.75s 3343.39MB 4109.18MB
Att-GRU 0.966 0.042 4247.81s 2.94s 3341.72MB 4111.50MB
ADTI
LSTM 0.930 0.416 303.00s 0.91s 995.50MB 932.71MB
GRU 0.933 0.411 270.90s 0.87s 966.58MB 946.81MB
BiGRU 0.935 0.403 288.48s 1.63s 3338.62MB 3531.10MB
Att-GRU 0.954 0.494 312.40s 1.34s 3345.10MB 3535.34MB
ETT
LSTM 0.906 0.292 17.39s 6E-3s 565.50MB 552.71MB
GRU 0.915 0.243 16.55s 4E-3s 554.11MB 540.46MB
BiGRU 0.844 0.373 19.39s 0.03s 934.01MB 954.30MB
Att-GRU 0.843 0.417 20.24s 0.04s 1047.29MB 1054.71MB
MIMIC
LSTM 0.703 0.377 1884.14s 98.92s 2371.65MB 2377.12MB
GRU 0.699 0.384 1168.72s 59.33s 1165.51MB 1168.10MB
BiGRU 0.681 0.366 1661.75s 84.92s 2265.17MB 2130.61MB
Att-GRU 0.598 0.376 1667.51s 71.34s 2390.36MB 3889.35MB
TABLE E5: Results of ablation experiments
Dataset
Explainer
SIC AUC Run time Memory usage
[…]
MNIST
Vanilla EBM 0.883 0.217 $\mathbf{1568.18}$s $\mathbf{1.02}$s $\mathbf{1354.19}$MB $\mathbf{1200.41}$MB
ADTI
Vanilla EBM 0.863 0.544 $\mathbf{173.22}$s $\mathbf{0.81}$s $\mathbf{889.21}$MB $\mathbf{850.41}$MB
EmSHAP $\mathbf{0.933}$ $\mathbf{0.411}$ 270.90s 0.87s 966.58MB 946.81MB
ETT
Vanilla EBM 0.864 0.440 $\mathbf{14.19}$s $\mathbf{1E-3}$s $\mathbf{515.72}$MB $\mathbf{510.33}$MB
Vanilla EBM+GRU couping 0.903 0.292 15.99s 4E-3s 553.37MB 540.10MB
EmSHAP $\mathbf{0.915}$ $\mathbf{0.243}$ 16.55s 4E-3s 554.11MB 540.46MB
MIMIC
Vanilla EBM 0.530 0.477 64.50s 31.21s 1051.62MB 1075.91MB
Vanilla EBM+Dynamic masking 0.560 0.464 69.25s 32.70s 1065.44MB 1079.08MB
Vanilla EBM+GRU coupling 0.667 0.386 1167.89s 59.33s 1074.56MB 1113.55MB
EmSHAP 0.699 0.384 1168.72s 59.93s 1165.51MB 1168.10MB